OpenStudy (anonymous):

could someone possibly help me out with a geometry question? I'm looking for an explanation as I've tried reading my lessons, as well as reaching out for help with other students of mine, and they haven't been able to help me. My teacher is away and I cannot call her right now -- anybody willing to help me with ONE geometry question? Pentagon PQRST and its reflection, pentagon P'Q'R'S'T', are shown in the coordinate plane below:

OpenStudy (anonymous):

What is the line of reflection between pentagons PQRST and P'Q'R'S'T'? x = 0 y = x y = 0 x = 1

OpenStudy (aihberkhan):

Okay. You just need to see the line in which the shape reflects. As you can see, the line reflects over the vertical line in the middle. What is that line called? Do you know? @osnapitzari

OpenStudy (jdoe0001):

ditto

OpenStudy (anonymous):

No, what is it? :)

OpenStudy (anonymous):

any information is helpful -- so I thank you in advance for helping me out with this :( I seem to be passing geometry with flying colors, but this seems to be the simplest, yet most difficult for me to comprehend :(

OpenStudy (jdoe0001):

the reflection occurs at the line AihberKhan indicated the y-axis, now, what is it called? well hint: \(\begin{array}{rrllll} x&y\\ 0&-\infty\\ 0&-1000\\ 0&-100\\ 0&0\\ 0&100\\ 0&1000\\ 0&1000000\\ 0&+\infty \end{array}\) what do you think is that line?

OpenStudy (anonymous):

is it called the mirror line?

OpenStudy (anonymous):

I remember something about that being mentioned

OpenStudy (jdoe0001):

well, yes hhmm but notice your choices :)

OpenStudy (jdoe0001):

notice your choices, and notice the values table \(\begin{array}{rrllll} x&y\\ 0&-\infty\\ 0&-1000\\ 0&-100\\ 0&0\\ 0&100\\ 0&1000\\ 0&1000000\\ 0&+\infty \end{array}\)

OpenStudy (anonymous):

oh...wOW, IT'S THAT EASY.

OpenStudy (jdoe0001):

no matter what value you plug into "y", "x" will always be 0 thus \(\large x= 0\) :)

OpenStudy (anonymous):

YOU'RE MAKING THIS SO CLEAR OMG

OpenStudy (anonymous):

Thank you so much! You gave me such a better understanding!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Is there any way to give you a medal -- or did I just do so by becoming a fan?

OpenStudy (jdoe0001):

tis ok :)